What is the electric potential 1 m away from a +1 nC charge? Group of answer choices 6 V 1 V 9 V 3 V
Answer #1

Solution:

Electric potential (V) = K (Q/r)

K = 9×10^9 N-m/C^2

Q= 1×10^-9 C

r = 1 m

V = 9×10^9 ×(1×10^-9 /1 )

V = 9 volts Answer